A solid gold vase of mass 20.0 kg is suspended by a rope at sea bottom. What is the Buoyant 

Force of the seawater on the vase?



1
Expert's answer
2021-07-12T12:17:32-0400

Solution.

"m=20.0kg;"

"\rho_g=19300kg/m^3;"

"\rho_w=1030kg/m^3;"

"F_B=\rho_wgV_g;"

"V_g=\dfrac{m}{\rho_g};"

"F_B=\dfrac{\rho_wgm}{\rho_g};"

"F_B=\dfrac{1030\sdot9.8\sdot20.0}{19300}=10.46N;"

Answer: "F_B=10.46N."